the electric flux through sphere of radiur 0.1m is 1.5vm. what is the flux through the sphere of radius .5m

Grade:12th pass

1 Answers

Divya
6 years ago
Electric flux us directly proportional to Area.Therefore Electric flux is directly proportional to radius.Hence,1.5/x=0.1/0.5x=1.5×0.5/0.1 = 7.5